Find all complex numbers that satisfy the equation 24/3 +8+8V3i = 0. Justify your answer!
Answer #1

De Moirres Formula (Goti Sino) = Go no + i sinne (6.0 +i Sino)* = G. 25W +O + i sin 2*** n where K=0,1,2, z-in 24/3 z +8+85/ч → Z = 24 = (24)37 Cos (0) +ísin(0) [124*{@(0)= isimco} } = (2)*+Go (4 mg) + (0+2k DO 0+ 2KT whene k=0,1,2,3 = 23 [ s Ces